OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

y=x+12

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

−x + y = 10 3x – 4y = −40 y = x + 12 y = x + 10

OpenStudy (anonymous):

Which answer choice is that ?

OpenStudy (anonymous):

3rd one

OpenStudy (anonymous):

Point slope form is \[y- y_{1} = m(x-x _{1})+b\] Given our equation \[y - 4 = (x +8)\] we see that our points are (-8, 4) Since we see no number in front of \[(x-x_1)\] we assume m =1 Now standard form is \[y = mx + b\] plugging in \[x=-8, y =4\] we find \[b = 12\] so our standard equation is \[y = x + 12\]